OpenStudy (anonymous):

how do you find the exact value of tan^-1(tan4pi/5)

OpenStudy (anonymous):

Let\[\theta =\tan [(4\pi)/5]\]Draw a triangle with theta. From the triangle find tan inverse of theta.

OpenStudy (anonymous):

wat happened to tan^-1

OpenStudy (anonymous):

Nothing happened. All problems have steps. From the triangle find tan^-1 of theta.

OpenStudy (anonymous):

i got pi/5 yet the answer (-pi/5)

OpenStudy (anonymous):

I don't know what you did.

OpenStudy (anonymous):

did u get (-pi/4)?

OpenStudy (anonymous):

sorry 5?

OpenStudy (anonymous):

I would think you would get some numbers representing sides of triangles.

OpenStudy (anonymous):

nah answer is (-pi/4)

OpenStudy (anonymous):

* (-pi/5)

OpenStudy (anonymous):

Don't concentrate on the answer. Use the answer as a check. Concentrate on the work.

OpenStudy (anonymous):

ye i cant seehow they got the answer thought

OpenStudy (anonymous):

Usually you are given a special angle to work with. 4pi/5 is not a special angle, I don't think, did you copy the question right?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

Go back to what I said earlier. Actually let 4pi/5=theta. Draw a triangle with theta as I labelled it. Call the side opposite theta x and the adj 1. Therefore tan of theta is x/1. It follows tan^-1 theta is 1/x. Theta is the angle whose tan is 1/x. I think they flipped it to 5pi/4, which after inspection should be same angle as -pi/5. Got to go, don't have time to check my thesis.

OpenStudy (anonymous):

k thanks

myininaya (myininaya):

tan inverse and tan are inverse functions they cancel each other and you should get 4pi/5 but how do we get -pi/5 let me think ...

myininaya (myininaya):

what is the restriction on the domain of tan inverse

myininaya (myininaya):

the domain is -pi/2 to pi/2 4pi/5 is definity not between this

myininaya (myininaya):

tan4pi/5=tan(-pi/5) 4pi/5 is over there in the cooridate (-,+) and -pi/5 is over ther ein the coodinate (+,-) and -pi/5 is in between -pi/2 and pi/2 so tan^(-1)(tan(-pi/5))=-pi/5

myininaya (myininaya):

and you there?

OpenStudy (anonymous):

thanks man